# ConCISE: Confidence-guided Compression in Step-by-step Efficient Reasoning

### v1

@@ B.3 Evaluation Dataset
Input Question: What day of the week will it be 40 days from Tuesday? Response: Step1: Alright, so I need to figure out what day of the week it will be 40 days from Tuesday. Hmm, okay. Let me think about how to approach this. Step2: First, I remember that each week has 7 days. So, if something happens every 7 days, it’s on the same day of the week next time. That might be useful here. Maybe I can figure out how many weeks are in 40 days and then see how many extra days that is. Step3: Let me do the division: 40 divided by 7. I think that’s 5 weeks and 5 days. Step4: Wait, let me check. 7 times 5 is 35, right? So, 40 minus 35 is 5. So, 40 days is equal to 5 weeks and 5 days. That makes sense. Step5: So, if today is a Tuesday, then 35 days later is also a Tuesday. Because 35 is exactly 5 weeks, which is 5 times 7. So, after 35 days, it’s still Tuesday. Now, I need to find out what day it is 5 days after Tuesday. Step6: Let me count: Tuesday plus one day is Wednesday, plus two days is Thursday, plus three days is Friday, plus four days is Saturday, and plus five days is Sunday. So, 5 days after Tuesday is $𝑆𝑢𝑛𝑑𝑎𝑦$. Step7: Therefore, 40 days from Tuesday is Sunday. Let me double-check that to make sure I didn’t make a mistake. If today is Tuesday and 7 days later is also Tuesday, then adding another 7 days each time cycles through the week. So, 35 days later, as we calculated, it is Tuesday again. Then, 5 days beyond that would be Sunday. Yeah, that seems right. Step8: Final Answer: The day of the week 40 days from Tuesday is $𝑆𝑢𝑛𝑑𝑎𝑦$.

### v2

@@ B.2 Training Details
Input Question: What day of the week will it be 40 days from Tuesday? Response: Step1: Alright, so I need to figure out what day of the week it will be 40 days from Tuesday. Hmm, okay. Let me think about how to approach this. Step2: First, I remember that each week has 7 days. So, if something happens every 7 days, it’s on the same day of the week next time. That might be useful here. Maybe I can figure out how many weeks are in 40 days and then see how many extra days that is. Step3: Let me do the division: 40 divided by 7. I think that’s 5 weeks and 5 days. Step4: Wait, let me check. 7 times 5 is 35, right? So, 40 minus 35 is 5. So, 40 days is equal to 5 weeks and 5 days. That makes sense. Step5: So, if today is a Tuesday, then 35 days later is also a Tuesday. Because 35 is exactly 5 weeks, which is 5 times 7. So, after 35 days, it’s still Tuesday. Now, I need to find out what day it is 5 days after Tuesday. Step6: Let me count: Tuesday plus one day is Wednesday, plus two days is Thursday, plus three days is Friday, plus four days is Saturday, and plus five days is Sunday. So, 5 days after Tuesday is Sunday. Step7: Therefore, 40 days from Tuesday is Sunday. Let me double-check that to make sure I didn’t make a mistake. If today is Tuesday and 7 days later is also Tuesday, then adding another 7 days each time cycles through the week. So, 35 days later, as we calculated, it is Tuesday again. Then, 5 days beyond that would be Sunday. Yeah, that seems right. Step8: Final Answer: The day of the week 40 days from Tuesday is Sunday.